# A Priori Uncertainty Quantification of Reacting Turbulence Closure Models using Bayesian Neural Networks

### v1

@@ Abstract
While many physics-based closure model forms have been posited for the sub-filter scale (SFS) in large eddy simulation (LES), vast amounts of data available from direct numerical simulations (DNS) create opportunities to leverage data-driven modeling techniques. Albeit flexible, data-driven models still depend on the dataset and the functional form of the model chosen. Increased adoption of such models requires reliable uncertainty estimates both in the data-informed and out-of-distribution regimes. In this work, we employ Bayesian neural networks (BNNs) to capture both epistemic and aleatoric uncertainties in a reacting flow model. In particular, we model the filtered progress variable scalar dissipation rate which plays a key role in the dynamics of turbulent premixed flames. We demonstrate that BNN models can provide unique insights about the structure of uncertainty of the data-driven closure models. We also propose a method for the incorporation of out-of-distribution information in a BNN. The efficacy of the model is demonstrated by a priori evaluation on a dataset consisting of a variety of flame conditions and fuels.
@@ 1 Introduction
Numerical simulations of turbulent reacting flows are usually associated with high computational cost due to the large range of spatiotemporal scales that need to be resolved pope2000turbulent . Concurrent with significant computational advances alexander2020exascale , several modeling approaches have since been devised to avoid resolving the smallest spatio-temporal scales, at the expense of the need for closure modeling echekki2010turbulent , poinsot2005theoretical , pitsch2006large . While several physics-based modeling strategies have emerged, they tend to make strict assumptions about the unresolved scales fiorina2005premixed , sagaut2005large . While physics models increasingly avoid these assumptions akram2022approximate , data-driven strategies have become more and more popular given their flexibility ihme2022combustion , data availability, and the development of scientific machine learning (SciML) frameworks karniadakis2021physics , kovachki2021neural , peherstorfer2016data , duraisamy2019turbulence , duraisamy2021perspectives .
@@
Multiple data-driven strategies have been shown to be accurate approximators of closure terms, such as artificial neural networks (ANN) vollant2017subgrid , barwey2021data , maulik2019subgrid , raman2019emerging , van2023energy , ihme2022combustion and convolutional neural networks (CNN) lapeyre2019training , nikolaou2019progress . Recent advances in generative modeling have paved the way for deconvolutional strategies that have successfully been applied to closure modeling bode2021using . While flexible, data-driven closure modeling strategies heavily depend on the model form and training procedure, both of which are typically optimized via a hyper-parameter search. In addition, the choice of the dataset itself can induce significant variability on the closure model hassanaly2023uniform . These choices can induce a significant uncertainty thereby motivating the need to equip these machine learning (ML) methods with reliable uncertainty estimates that can be propagated through the governing equations.
@@
A variety of machine learning methods have been developed to model uncertainty hullermeier2021aleatoric , abdar2021review , nelsen2021random . Popular methods to quantify uncertainty in neural networks include the dropout method srivastava2014dropout , jospin2022hands and its Bayesian interpretation gal2016dropout . While inexpensive and easy to implement, dropout generates probabilistic predictions at inference time but does not quantify the model parameter uncertainty anderson2020meaningful . In computational science and engineering, Gaussian processes have long been applied to uncertainty quantification tasks rasmussen2006gaussian owing to their flexibility and interpretability. The primary drawback of Gaussian processes is the $\mathcal{O}(n^{3})$ computation expense to train and $\mathcal{O}(n^{2})$ cost to evaluate due to the required matrix inversion and multiplication, where $n$ is the data size. Despite efforts, scalability for large datasets remains a challenge snelson2008flexible , liu2020gaussian , banerjee2013efficient , wang2016bayesian , bauer2016understanding . New approaches are needed to handle massive datasets generated by DNS, especially in the context of rapid online evaluation within LES codes.
@@
BNNs are an attractive method to estimate and predict modeling uncertainty due to their ability to ingest large amounts of data, relatively fast inference cost (compared to Gaussian processes), rigorous characterization of uncertainty, and expressivity. Recent advancements employing variational inference jospin2022hands have made the training of BNNs tractable for large models and amenable to large datasets graves2011practical , blundell2015weight . BNNs reformulate deterministic deep learning models as point-estimators and emulate the construction of an ensemble of neural nets by assigning a probability distribution to each network parameter mackay1995probable . Thus, they generate a predictive distribution by sampling the parameter distributions and collecting the resulting distribution of point estimates. Recent applications of BNNs for combustion modeling have mostly focused on the prediction of macroscopic quantities such as ignition delay mccartney2022reducing , fuel properties oh2024learning , or to accelerate data assimilation croci2021data . This work is the first to explore the use of BNNs for quantifying both epistemic and aleatoric uncertainties stemming from the adopted model form and training dataset najm2009uncertainty . As an illustrative exercise, the focus of this work is on applying the BNN methodology for modeling sub-filter progress variable dissipation rate of premixed turbulent flames yellapantula2021deep . However, the framework developed in this work could apply to any other data-driven closure modeling type.
@@ 2.1 Sources of Uncertainty
The nature of uncertainty in numerical simulations and how to deal with their influence is of particular importance in engineering contexts where one optimizes designs according to specific quantities of interest, e.g. ignition time tang2021probabilistic , jaravel2019numerical or maximum temperature masquelet2017uncertainty . While there are multiple sources of uncertainty arising from model inputs, numerical errors, or experiments, we primarily consider the uncertainties driven by the data that the model trains on. In the machine learning context, the use of deep neural networks is typically justified by a universal approximation theorem hornik1989multilayer valid in the asymptotic limit of data availability. Thus, we concern ourselves with uncertainty in the learned parameters and resultant functional representation due to the empirical data distribution used for training.
Broadly speaking, uncertainties can be categorized as either aleatoric or epistemic smith2013uncertainty , der2009aleatory , najm2009uncertainty . As an illustrative example, consider the example shown in Fig. 1. The underlying generating function for the data is
@@
In the context of data-driven closure modeling, epistemic uncertainties arise from a lack of knowledge when too few training data are available. An ideal LES model in the sense of Langford et al. langford1999optimal obtained via a data-driven method, could be subject to epistemic uncertainty when there does not exist sufficient data to confidently determine the optimal estimator of the closure term. Epistemic uncertainty also presents itself in the extrapolation setting, which may be viewed as an extreme case of lacking data availability. In addition, since epistemic uncertainty may be mitigated with additional data, it can guide where additional data is needed.
Aleatoric uncertainty is also known as statistical or stochastic uncertainty and is inherent to a problem or experiment. It is irreducible in the sense that it cannot be reduced by additional data collection. A clear example is the measurement tolerance of a sensor: additional data collection would not inform one to a greater precision than the instrument rating. In the data-driven closure modeling context, aleatoric uncertainty primarily presents itself as coarse–graining or filtering uncertainty. The act of filtering destroys information, as there are potentially many DNS realizations that, when filtered, would present identical filtered measurements. Inferring a mapping from filtered to unfiltered data is therefore subject to an irreducible uncertainty langford1999optimal , adrian1990stochastic , pope2010self , hassanaly2022adversarial . Likewise, the selection of input features can also be seen as a coarse-graining step. Unlike epistemic uncertainty, aleatoric uncertainty does not guide the acquisition of data but rather can guide further refinement in the choice of the input features by highlighting regions of phase space where the model may be unreliable.
@@ 2.2 Reacting Turbulent Flow Dataset Generation
The dataset used in this work is obtained from several DNS of turbulent premixed flames lapointe2015differential , savard2017effects . The reader is referred to Yellapantula et al. yellapantula2021deep for a more detailed description of the data generation and preparation processes. Hereafter, only key details of the dataset are repeated. The DNS data was Favre filtered using Gaussian filter kernels with different filter sizes to prepare input features for a priori model evaluation that would match the LES features available in the a posteriori setting. The objective is to learn an optimal mapping between the filtered quantities and the quantity of interest, here the unresolved contribution to the filtered progress variable dissipation rate, $\chi_{\rm SFS}$ defined as
@@
The relevance of approximating closure models for the subfilter progress variable dissipation rate has been extensively documented elsewhere and is not repeated here gao2014scalar , yellapantula2021deep .
The input features are the same as those considered in Ref. yellapantula2021deep and are summarized in Table 1. The input features were selected based on a feature importance metric sundararajan2017axiomatic and only contain filtered quantities as the model is envisioned to be deployed online within a larger LES model. Unlike Ref. yellapantula2021deep , the output layer only predicts the unresolved contribution of the progress variable dissipation rate $\chi_{\rm SFS}$. Other derived quantities include the principle rates of filtered strain rate. The principal rates of strain, $\alpha>\beta>\gamma$ are computed using the strain rate tensor constructed from the filtered velocity field. Additionally, the alignments between the eigenvectors and local gradient of the filtered progress variable were added to the list of input features yellapantula2021deep . The input space $\mathbf{x}\in\mathbb{R}^{d_{i}}$ is mapped to an output $\mathbf{y}\in\mathbb{R}^{d_{o}}$, where $d_{i}=10$ is the dimensionality of the input space and $d_{o}$ is the dimensionality of the output space, in this case $d_{o}=1$. The effect of the number of input parameters on the estimated uncertainties is further discussed in A.
@@
After the selection of the input features, a near-uniform in phase-space sampling hassanaly2023uniform was performed by clustering the data with 40 clusters in the input feature space and uniformly selecting data from each cluster. This process allows for appropriately capturing the reaction zone of all the DNS cases in the dataset with a reduced number of data points. The final dataset contains $7.88\times 10^{6}$ data points for training and $2.63\times 10^{6}$ data points for testing, which is about $10^{3}$ times less than the total number of DNS data points initially available.
@@ 2.3 Bayesian Neural Network Modeling
Classical neural networks are only capable of deterministically generating point estimates and thus are not well suited for the task of assessing uncertainty. On the other hand, Bayesian neural networks model epistemic uncertainty estimation by placing a parametric distribution over the neural network parameters denker1990transforming , mackay1992practical , neal2012bayesian , blundell2015weight . By sampling the parameter distributions, one selects an ensemble of deterministic neural networks. Note that sampling multiple candidate neural networks to estimate uncertainty is consistent with peer-modeling approaches for uncertainty propagation mueller2018model . In particular, we consider the parametric distribution for the weights to be a Gaussian blundell2015weight . The mean and variance of which become the parameters to be learned in the BNN representation. To delineate between the sampled parameters defining a deterministic network and the distribution placed over the weights in the BNN representation, we will set the notation that “weights” $\mathbf{w}$ are sampled from a (parametric) distribution placed upon the weights $q(\mathbf{w}|\theta)$, where $\theta$ are the parameters of the BNN. Under this notation, $\mathbf{w}\sim q(\mathbf{w}|\theta)$.
In the variational inference formulation of BNN training magris2023bayesian , one seeks to minimize the Kullback-Liebler (KL) divergence between the weight distribution and the true Bayesian posterior conditioned on the dataset $\boldsymbol{D}$,
@@
Expansion and rearrangement of this formulation yields the evidence lower bound (ELBO) objective function blundell2015weight ,
@@
Although the primary interest of modelers that deploy closure models in a reacting flow solver is the epistemic uncertainty, one may wish to evaluate the aleatoric uncertainty as well (see Sec. 2). One method for doing so is to further parameterize each output dimension so as to capture heteroskedastic uncertainty in the data kendall2017uncertainties , hoffmann2021deep , nemani2023uncertainty . We employ the modeling choice from Kendall and Gal kendall2017uncertainties of a Gaussian form for the output random variable.
@@
Here $x_{\ell}$ denotes the output from the last layer, and $\mathbf{x}\in\mathbb{R}^{d_{i}}$ again represents an input data point. The final layer is mapped to $(\mathbf{y}_{\boldsymbol{\mu}},\mathbf{y}_{\boldsymbol{\sigma}})^{T}\in\mathbb{R}^{2d_{o}}$ in the case of a Kendall and Gal kendall2017uncertainties style architecture. Here $\mathbf{y}_{\boldsymbol{\mu}},\mathbf{y}_{\boldsymbol{\sigma}}$ parameterize the output variable as $\mathbf{y}\in\mathbb{R}^{d_{o}}\sim\mathcal{N}\left(\mathbf{y}_{\boldsymbol{\mu}},\operatorname{diag}(\mathbf{y}_{\boldsymbol{\sigma}})\right)$. We prefer this parameterization, because if we were to directly map to the output dimension, $\mathbf{y}\in\mathbb{R}^{d_{o}}$ and directly predict the output quantity, only the epistemic uncertainty would be modeled. The difference in the two architectures is shown graphically for $d_{o}=1$ in Fig. 1(a) and Fig. 1(b), respectively. In this work, we adopt the notation that the output $\mathbf{y}$ is the result of sampling the output of a BNN parameterized by $\theta$ at a datum $\mathbf{x}$ after appropriate sampling of the weights $\mathbf{w}$ and the resultant random variable.
In this framework, one can use the law of total variance saltelli2010variance to decompose the variance (uncertainty) in the predictions into its epistemic and aleatoric components chai2018uncertainty , nemani2023uncertainty ,
$\operatorname{Var}(\mathbf{y})=\underbrace{\mathbb{E}_{q(\mathbf{w}|\theta)}\left[\operatorname{Var}(\mathbf{y}|\mathbf{x},\boldsymbol{D})\right]}_{\text{aleatoric}}+\underbrace{\operatorname{Var}\left(\mathbb{E}_{q(\mathbf{w}|\theta)}\left[\mathbf{y}|\mathbf{x},\boldsymbol{D}\right]\right)}_{\text{epistemic}}.$ (8)
The predictive variance $\operatorname{Var}(\mathbf{y})$ is decomposed into an aleatoric component $\mathbb{E}_{q(\mathbf{w}|\theta)}\left[\operatorname{Var}(\mathbf{y}|\mathbf{x},\boldsymbol{D})\right]$, the mean variability in the estimates, and an epistemic component $\operatorname{Var}(\mathbb{E}_{q(\mathbf{w}|\theta)}[\mathbf{y}|\mathbf{x},\boldsymbol{D}])$, the variability in the model’s mean predictions. For a given $\mathbf{w}\sim q(\mathbf{w}|\theta)$, the predictive mean is given by the first network output, $\mathbf{y}_{\boldsymbol{\mu}}\in\mathbb{R}^{d_{o}}$. Detailed algorithms for computation of the uncertainties are provided in B. For a concrete example, the magnitudes of the uncertainties for the one-dimensional example (Fig. 1) are shown in Fig. 3. It can be observed that the regions with high and low epistemic uncertainty are appropriately characterized in Fig. 2(a) and that the predictive envelope captures the entire data distribution in Fig. 2(b).
Figure 2: (a) A BNN in the style of blundell2015weight that captures epistemic uncertainty only and (b) a BNN in the style of kendall2017uncertainties that captures both epistemic and aleatoric uncertainty.

### v2

@@ Abstract
While many physics-based closure model forms have been posited for the sub-filter scale (SFS) in large eddy simulation (LES), vast amounts of data available from direct numerical simulations (DNS) create opportunities to leverage data-driven modeling techniques. Albeit flexible, data-driven models still depend on the dataset and the functional form of the model chosen. Increased adoption of such models requires reliable uncertainty estimates both in the data-informed and out-of-distribution regimes. In this work, we employ Bayesian neural networks (BNNs) to capture both epistemic and aleatoric uncertainties in a reacting flow model. In particular, we model the filtered progress variable scalar dissipation rate which plays a key role in the dynamics of turbulent premixed flames. We demonstrate that BNN models can provide unique insights about the structure of uncertainty of the data-driven closure models. We also propose a method for the incorporation of out-of-distribution information in a BNN, which can be used for out-of-distribution query detection. The efficacy of the model is demonstrated by a priori evaluation on a dataset consisting of a variety of flame conditions and fuels.
@@ 1 Introduction
Numerical simulations of turbulent reacting flows are usually associated with high computational cost due to the large range of spatiotemporal scales that need to be resolved [1]. Concurrent with significant computational advances [2], several modeling approaches have since been devised to avoid resolving the smallest spatio-temporal scales, at the expense of the need for closure modeling [3, 4, 5]. While several physics-based modeling strategies have emerged, they tend to make strict assumptions about the unresolved scales [6, 7]. While physics models increasingly avoid these assumptions [8], data-driven strategies have become more and more popular given their flexibility [9], data availability, and the development of scientific machine learning (SciML) frameworks [10, 11, 12, 13, 14, 15].
@@
There exists a large body of literature deploying SciML techniques for closure modeling in general [16]. For turbulent combustion applications, multiple data-driven strategies have been shown to be accurate approximators of closure terms, such as artificial neural networks [17, 18, 19, 20, 21, 9] and convolutional neural networks [22, 23]. Recent advances in generative modeling have led to successful deconvolution strategies applied to closure modeling [24]. While flexible, data-driven closure modeling strategies still depend on the model form and the training procedure, both of which are typically optimized via a hyper-parameter search. In addition, the choice of the dataset itself can induce significant variability on the closure model [25]. These choices can result in a significant uncertainty thereby motivating the need to equip these machine learning (ML) methods with reliable uncertainty estimates that can be propagated through the governing equations.
@@
A variety of machine learning methods have been developed to model uncertainty [26, 27, 28]. Popular methods to quantify uncertainty in neural networks include the dropout method [29, 30] and its Bayesian interpretation [31]. While inexpensive and easy to implement, dropout generates probabilistic predictions at inference time but does not quantify the model parameter uncertainty [32]. In computational science and engineering, Gaussian processes have long been applied to uncertainty quantification tasks [33] owing to their flexibility and interpretability. The primary drawback of Gaussian processes is the $\mathcal{O}(n^{3})$ computation expense to train and $\mathcal{O}(n^{2})$ cost to evaluate due to the required matrix inversion and multiplication, where $n$ is the data size. Despite efforts, scalability for large datasets remains a challenge [34, 35, 36, 37, 38]. New approaches are needed to handle massive datasets generated by DNS, especially in the context of rapid online evaluation within LES codes.
@@
BNNs are an attractive method to estimate and predict modeling uncertainty due to their ability to ingest large amounts of data, relatively fast inference cost (compared to Gaussian processes), rigorous characterization of uncertainty, and expressivity. Recent advancements employing variational inference [30] have made the training of BNNs tractable for large models and amenable to large datasets [39, 40]. BNNs reformulate deterministic deep learning models as point-estimators and emulate the construction of an ensemble of neural nets by assigning a probability distribution to each network parameter [41]. Thus, they generate a predictive distribution by sampling the parameter distributions and collecting the resulting distribution of point estimates. Recent applications of BNNs for combustion modeling have mostly focused on the prediction of macroscopic quantities such as ignition delay [42], fuel properties [43], or to accelerate data assimilation [44]. This work is the first to explore the use of BNNs for quantifying both epistemic and aleatoric uncertainties stemming from the adopted model form and training dataset [45]. As an illustrative exercise, the focus of this work is on applying the BNN methodology for modeling sub-filter progress variable dissipation rate of premixed turbulent flames [46]. However, the framework developed in this work could apply to any other data-driven closure modeling type.
@@ 2.1 Sources of Uncertainty
The nature of uncertainty in numerical simulations and how to deal with their influence is of particular importance in engineering contexts where one optimizes designs according to specific quantities of interest, e.g. ignition time [47, 48] or maximum temperature [49]. While there are multiple sources of uncertainty arising from model inputs, numerical errors, or experiments, we primarily consider the uncertainties driven by the data that the model trains on. In the machine learning context, the use of deep neural networks is typically justified by a universal approximation theorem [50] valid in the asymptotic limit of data availability. Thus, we concern ourselves with uncertainty in the learned parameters and resultant functional representation due to the empirical data distribution used for training.
Broadly speaking, uncertainties can be categorized as either aleatoric or epistemic [51, 52, 45]. As an illustrative example, consider the example shown in Fig. 1. The underlying generating function for the data is
@@
In the context of data-driven closure modeling, epistemic uncertainties arise from a lack of knowledge when too few training data are available. An ideal LES model in the sense of Langford et al. [53] obtained via a data-driven method, could be subject to epistemic uncertainty when there does not exist sufficient data to confidently determine the optimal estimator of the closure term. Epistemic uncertainty also presents itself in the extrapolation setting, which may be viewed as an extreme case of lacking data availability. In addition, since epistemic uncertainty may be mitigated with additional data, it can guide where additional data is needed.
Aleatoric uncertainty is also known as statistical or stochastic uncertainty and is inherent to a problem or experiment. It is irreducible in the sense that it cannot be reduced by additional data collection. A clear example is the measurement tolerance of a sensor: additional data collection would not inform one to a greater precision than the instrument rating. In the data-driven closure modeling context, aleatoric uncertainty primarily presents itself as coarse–graining or filtering uncertainty. The act of filtering destroys information, as there are potentially many DNS realizations that, when filtered, would present identical filtered measurements. Inferring a mapping from filtered to unfiltered data is therefore subject to an irreducible uncertainty [53, 54, 55, 56]. Likewise, the selection of input features can also be seen as a coarse-graining step. Unlike epistemic uncertainty, aleatoric uncertainty does not guide the acquisition of data but rather can guide further refinement in the choice of the input features by highlighting regions of phase space where the model may be unreliable.
@@ 2.2 Reacting Turbulent Flow Dataset Generation
The dataset used in this work is obtained from several DNS of turbulent premixed flames [57, 58]. The reader is referred to Yellapantula et al. [46] for a more detailed description of the data generation and preparation processes. Hereafter, only key details of the dataset are repeated. The DNS data was Favre filtered using Gaussian filter kernels with different filter sizes to prepare input features for a priori model evaluation that would match the LES features available in the a posteriori setting. The objective is to learn an optimal mapping between the filtered quantities and the quantity of interest, here the unresolved contribution to the filtered progress variable dissipation rate, $\chi_{\rm SFS}$ defined as
@@
The relevance of approximating closure models for the subfilter progress variable dissipation rate has been extensively documented elsewhere and is not repeated here [59, 46].
The input features are the same as those considered in Ref. [46] and are summarized in Table 1. The input features were selected based on a feature importance metric [60] and only contain filtered quantities as the model is envisioned to be deployed online within a larger LES model. Unlike Ref. [46], the output layer only predicts the unresolved contribution of the progress variable dissipation rate $\chi_{\rm SFS}$. Other derived quantities include the principle rates of filtered strain rate. The principal rates of strain, $\alpha>\beta>\gamma$ are computed using the strain rate tensor constructed from the filtered velocity field. Additionally, the alignments between the eigenvectors and local gradient of the filtered progress variable were added to the list of input features [46]. The input space $\mathbf{x}\in\mathbb{R}^{d_{i}}$ is mapped to an output $\mathbf{y}\in\mathbb{R}^{d_{o}}$, where $d_{i}=10$ is the dimensionality of the input space and $d_{o}$ is the dimensionality of the output space, in this case $d_{o}=1$. The effect of the number of input parameters on the estimated uncertainties is further discussed in A.
@@
After the selection of the input features, a near-uniform in phase-space sampling [25] was performed by clustering the data with 40 clusters in the input feature space and uniformly selecting data from each cluster. This process allows for appropriately capturing the reaction zone of all the DNS cases in the dataset with a reduced number of data points. The final dataset contains $7.88\times 10^{6}$ data points for training and $2.63\times 10^{6}$ data points for testing, which is about $10^{3}$ times less than the total number of DNS data points initially available.
@@ 2.3 Bayesian Neural Network Modeling
Classical neural networks are only capable of deterministically generating point estimates and thus are not well suited for the task of assessing uncertainty. On the other hand, Bayesian neural networks model epistemic uncertainty estimation by placing a parametric distribution over the neural network parameters [61, 62, 63, 40]. By sampling the parameter distributions, one selects an ensemble of deterministic neural networks. Note that sampling multiple candidate neural networks to estimate uncertainty is consistent with peer-modeling approaches for uncertainty propagation [64]. In particular, we consider the parametric distribution for the weights to be a Gaussian [40]. The mean and variance of the weights become the parameters to be learned in the BNN representation. To delineate between the sampled parameters defining a deterministic network and the distribution placed over the weights in the BNN representation, we will set the notation that “weights” $\mathbf{w}$ are sampled from a (parametric) distribution placed upon the weights $q(\mathbf{w}|\theta)$, where $\theta$ are the parameters of the BNN. Under this notation, $\mathbf{w}\sim q(\mathbf{w}|\theta)$.
In the variational inference formulation of BNN training [65], one seeks to minimize the Kullback-Liebler (KL) divergence between the weight distribution and the true Bayesian posterior conditioned on the dataset $\boldsymbol{D}$,
@@
Expansion and rearrangement of this formulation yields the evidence lower bound (ELBO) objective function [40],
@@
Although the primary interest of modelers that deploy closure models in a reacting flow solver is the epistemic uncertainty, one may wish to evaluate the aleatoric uncertainty as well (see Sec. 2). One method for doing so is to further parameterize each output dimension so as to capture heteroskedastic uncertainty in the data [66, 67, 68]. We employ the modeling choice from Kendall and Gal [66] of a Gaussian form for the output random variable.
@@
Here $x_{\ell}$ denotes the output from the last layer, and $\mathbf{x}\in\mathbb{R}^{d_{i}}$ again represents an input data point. The final layer is mapped to $(\mathbf{y}_{\boldsymbol{\mu}},\mathbf{y}_{\boldsymbol{\sigma}})^{T}\in\mathbb{R}^{2d_{o}}$ in the case of a Kendall and Gal [66] style architecture. Here $\mathbf{y}_{\boldsymbol{\mu}},\mathbf{y}_{\boldsymbol{\sigma}}$ parameterize the distribution of the output variable as $\mathbf{y}\in\mathbb{R}^{d_{o}}\sim\mathcal{N}\left(\mathbf{y}_{\boldsymbol{\mu}},\operatorname{diag}(\mathbf{y}_{\boldsymbol{\sigma}})\right)$. This parameterization allow to model epistemic uncertainty (through the variance of $\mathbf{y}_{\boldsymbol{\mu}}$) the aleatoric uncertainty (through the variance of $\mathbf{y}_{\boldsymbol{\sigma}}$). The difference in the epistemic only architecture and the epistemic and aleatoric architectures is shown graphically for $d_{o}=1$ in Fig. 2(a) and Fig. 2(b), respectively. In this work, we adopt the notation that the output $\mathbf{y}$ is the result of sampling the output of a BNN parameterized by $\theta$ at a datum $\mathbf{x}$ after appropriate sampling of the weights $\mathbf{w}$ and the resultant random variable.
In this framework, one can use the law of total variance [69] to decompose the variance (uncertainty) in the predictions into its epistemic and aleatoric components [70, 68],
$\operatorname{Var}(\mathbf{y})=\underbrace{\mathbb{E}_{q(\mathbf{w}|\theta)}\left[\operatorname{Var}(\mathbf{y}|\mathbf{x})\right]}_{\text{aleatoric}}+\underbrace{\operatorname{Var}_{q(\mathbf{w}|\theta)}\left(\mathbb{E}\left[\mathbf{y}|\mathbf{x}\right]\right)}_{\text{epistemic}}.$ (8)
The predictive variance $\operatorname{Var}(\mathbf{y})$ is decomposed into an aleatoric component $\mathbb{E}_{q(\mathbf{w}|\theta)}\left[\operatorname{Var}(\mathbf{y}|\mathbf{x})\right]$, the mean variability in the estimates, and an epistemic component $\operatorname{Var}_{q(\mathbf{w}|\theta)}(\mathbb{E}[\mathbf{y}|\mathbf{x}])$, the variability in the model’s mean predictions. For a given $\mathbf{w}\sim q(\mathbf{w}|\theta)$, the predictive mean is given by the first network output, $\mathbf{y}_{\boldsymbol{\mu}}\in\mathbb{R}^{d_{o}}$. Detailed algorithms for computation of the uncertainties are provided in B. For a concrete example, the magnitudes of the uncertainties for the one-dimensional example (Fig. 1) are shown in Fig. 3. It can be observed that the regions with high and low epistemic uncertainty are appropriately characterized in Fig. 3(a) and that the predictive envelope captures the entire data distribution in Fig. 3(b).
Figure 2: (a) A BNN in the style of [40] that captures epistemic uncertainty only and (b) a BNN in the style of [66] that captures both epistemic and aleatoric uncertainty.